# Optimizing Task Scheduling in Fog Computing with Deadline Awareness

### v3

@@ I Introduction
In recent years, the Internet of Things (IoT) has become a prominent technology in the internet and networking industry [1]. It consists of a system of linked physical devices, like household items and cars, that have sensors, software, and online connectivity [14]. Although IoT has been expanding quickly, it is encountering obstacles such as traffic congestion, delays, and excessive energy usage.
Adhering to deadlines is crucial in real-time IoT applications such as: e-health in healthcare, smart grid in energy management, livestock monitoring in agriculture, and smart traffic in transportation management. These time-sensitive applications have stringent QoS requirements including response time, deadline adherence, and energy efficiency that exceed the capabilities of resource-constrained IoT devices [19]. Managing resources while minimizing energy usage and meeting deadlines remains challenging [20].
@@
Fog computing addresses these challenges by providing computational capabilities at the edge, reducing delays and energy consumption [18, 15]. However, fog computing’s diverse and limited resources necessitate efficient scheduling and allocation strategies [3]. Many existing solutions rely on single optimization strategies or fail to account for the dynamic, heterogeneous nature of IoT-Fog environments [21, 8, 2, 6].
@@
RIGEO operates within a three-layered architecture: cloud, fog, and IoT devices. The fog layer comprises a controller and a network for resource management and task scheduling, while the cloud offers significant computing and storage capabilities (Figure 1). Since task scheduling is NP-hard [10, 17], machine learning and heuristic methods are essential to achieve optimal performance [4].
Figure 1: The architecture of the IoT-Fog-Cloud network.
Task scheduling faces multiple challenges, including cost optimization, energy efficiency, meeting deadline constraints, reducing latency, ensuring reliability, and maintaining fault tolerance. The following subsections review relevant studies in these areas.
@@ I-A Energy Efficiency
The authors in [25] proposed MEETS, a new task scheduling algorithm for fog computing environments. The algorithm employs a two-step approach: determining the optimal task allocation to fog nodes, followed by scheduling tasks on each node. The results show that MEETS significantly reduces energy consumption while meeting all task deadlines. However, it assumes homogeneous fog networks and constant task arrival rates, which may not be realistic in practice.
The authors in [23] proposed Energy-Efficient Task Scheduling in Fog Computing based on Particle Swarm Optimization (EETSPSO). EETSPSO outperforms existing approaches in terms of makespan, energy consumption, and execution time. However, as fog nodes are resource-constrained, the algorithm’s performance may vary across different real-world scenarios with varying resource limitations.
@@ I-B Deadline Constraints
The task scheduling approach in [3] incorporates two semi-greedy algorithms: Priority-aware Semi-Greedy (PSG) and PSG with Multi-start Procedure (PSG-M). These algorithms optimize energy consumption and minimize deadline violations for IoT tasks by effectively mapping them to fog nodes while ensuring QoS compliance. The PSG-M algorithm demonstrates a significant reduction in deadline violation time, however, the performance may vary under different real-world scenarios.
In [12], Louail et al. introduced a dynamic task scheduling algorithm considering both deadline constraints and task frequency to improve fog node efficiency in smart factories. By considering task frequency, the algorithm optimizes resource utilization, reducing idle time and improving overall efficiency. However, the complexity and computational requirements of the algorithm could pose a challenge in resource-constrained environments.
@@ I-C Latency Reduction
Authors in [11] proposed an Energy-Efficient Internet of Medical Things (EEIoMT) to Fog Interoperability of Task Scheduling framework. This framework schedules tasks efficiently by ensuring that critical tasks are executed within deadlines while balancing energy consumption. By utilizing ECG sensors to monitor heart health in smart cities, the framework processes data closer to end-users, thereby reducing latency and enhancing response times for IoMT applications. However, the study focuses on specific scenarios, and performance may vary in different real-world conditions.
The approach in [9] presents the Priority Queue Fuzzy Analytical Hierarchy Process (PQFAHP) algorithm. Using fuzzy logic and AHP, it prioritizes tasks based on completion time, energy consumption, RAM usage, and deadlines. The goal is to achieve optimal trade-offs between cost and latency in mobile fog computing environments. However, implementation complexity and computational requirements could challenge resource-constrained environments.
I-D Cost Optimization
The authors in [21] proposed a Multi-Objective Cost-Aware Discrete Grey Wolf Optimization-based Algorithm (MoDGWA) to tackle task scheduling challenges in IoT applications. This algorithm optimizes execution time, costs, and reliability by incorporating both execution and monetary cost models. Despite effectiveness, the algorithm’s complexity and computational demands could pose challenges in limited-resource environments.
The authors in [16] introduced a Cost-Aware Genetic-Based (CAG) task scheduling algorithm for fog-cloud environments. CAG improves cost efficiency in real-time applications with stringent deadlines by optimizing service and resource allocation within the three-tier IoT architecture. While effective for delay-sensitive applications, the study’s limited scenarios and metrics leave room for performance variability under different real-world conditions.
I-E Reliability and Fault Tolerance
In [8], the authors propose Dynamic Fault Tolerant Learning Automata (DFTLA) task scheduling, utilizing variable-structure learning automata to efficiently assign tasks to fog nodes, thereby ensuring reliable execution while optimizing response time and energy consumption. The authors in [2] proposed a Load Balanced Service Scheduling Approach (LBSSA) for cloud-fog environments, categorizing requests into real-time, important, and time-tolerant groups while incorporating resource failure rates to enhance reliability. However, its focus on specific scenarios may limit its broader applicability.
@@ I-E Reliability and Fault Tolerance
Energy Deadline Latency Cost Reliability
Eff. Const. Red. Opt.
[25] 2018 × × ×
[23] 2023 × × × ×
@@ II-A Problem Statement
The problem is to find an optimal mapping $M:T\rightarrow F$ to achieve these objectives. The parameters used in the methodology are defined as follows:
@@ II-A3 Energy Consumption
II-B Proposed Method (RIGEO)
@@
II-B1 Reinforcement Learning
RL is one of the main types of Machine Learning alongside Supervised, Unsupervised, and Semi-supervised Learning. RL algorithms train agents to make optimal decisions in complex environments by providing feedback and observing outcomes. The agent interacts with the environment, receives reinforcement signals, and learns to maximize long-term rewards by selecting optimal actions, enhancing decision-making capabilities through experience. The model comprises environmental states, agent actions, reinforcement signals, and an input function determining agent perception.
Initially, we create an array with a length equal to the long-deadline task count, where each cell references a task and its assigned FN. FNs can accommodate multiple tasks, with initial assignments randomly determined. This array evolves based on environmental feedback: RL evaluates the configuration, returning rewards when state $T+1$ improves upon state $T$, and penalties otherwise. Through this feedback mechanism, optimal scheduling is determined.
II-B2 Golden Eagle Optimization (GEO)
@@
II-B3 Improved Golden Eagle Optimization (IGEO)
@@
Negative step vectors in Eq. (7) indicate exploration tendency; positive vectors indicate exploitation. IGEO employs different genetic operators accordingly. For negative step vectors, mutation replaces addition in Eq. (8), randomly altering the parental genetic material as shown in Figure 2.
@@
For positive step vectors, crossover operators provide efficiency. Genes from two parents exchange at predetermined cutting points, shown in Figure 3. Eq. (10) shows the update for positive operations.
@@ III Results
Experimental results demonstrate that RIGEO reduces response time, minimizes deadline violation time, and optimizes energy consumption compared to other algorithms. The following sections examine RIGEO’s performance on each parameter.
@@ III-A Energy Consumption
Figure 4 illustrates the energy usage of FNs with varying task numbers (200 to 600) and 20 FNs. Energy consumption is calculated using Eq. (5).
@@
Figure 4: Average Total Energy Consumption (Joules) with varying task numbers (200-600) and 20 FNs over 50 runs
@@ III-B Deadline Violation
RIGEO improves deadline adherence compared to other methods, as indicated by the decrease in violation times shown in Figure 5. The method consistently selects nodes with the shortest violation times, even in the presence of node shortages, demonstrating flexibility in addressing deadline breaches.
@@
Figure 5: Average Deadline Violation Time (ms) with varying task numbers (200-600) and 20 FNs over 50 runs
@@ III-C Response Time
Response time, the period required for the system response to a task, was systematically integrated into the fitness function as a critical parameter. Figure 6 illustrates RIGEO’s superior performance in minimizing response time compared to other approaches with 20 FNs.
@@
Figure 6: Average Total Response Time (ms) with varying task numbers (200-600) and 20 FNs over 50 runs

### v4

@@ I Introduction
In recent years, the Internet of Things (IoT) has become a prominent technology in the internet and networking industry. It consists of a system of linked physical devices, like household items and cars, that have sensors, software, and online connectivity [9]. Although IoT has been expanding quickly, it is encountering obstacles such as traffic congestion, delays, and excessive energy usage.
Adhering to deadlines is crucial in real-time IoT applications such as: e-health in healthcare, smart grid in energy management, livestock monitoring in agriculture, and smart traffic in transportation management. These time-sensitive applications have stringent QoS requirements including response time, deadline adherence, and energy efficiency that exceed the capabilities of resource-constrained IoT devices [3].
@@
Fog computing addresses these challenges by providing computational capabilities at the edge, reducing delays and energy consumption. However, fog computing’s diverse and limited resources necessitate efficient scheduling and allocation strategies [1]. Many existing solutions rely on single optimization strategies or fail to account for the dynamic, heterogeneous nature of IoT-Fog environments [5].
Figure 1: The architecture of the IoT-Fog-Cloud network.
@@
RIGEO operates within a three-layered architecture: cloud, fog, and IoT devices. Figure 1 illustrates this hierarchy, where the fog layer comprises a controller and a network for resource management and task scheduling, while the cloud offers significant computing and storage capabilities. Since task scheduling is NP-hard [10], machine learning and heuristic methods are essential to achieve optimal performance [11].
The feasibility advantage of RIGEO over approaches that apply a single meta-heuristic to all tasks stems from matching algorithmic cost to task urgency. Population-based optimizers such as IGEO evaluate $N_{pop}\times N_{iter}$ candidate solutions per scheduling round, an overhead that is justified for short-deadline tasks, where even marginal improvements directly prevent costly deadline violations. For long-deadline tasks, however, this computational investment is excessive: a lightweight RL policy selects fog node assignments in $\mathcal{O}(|F_{\text{high}}|)$ per task by following a continuously updated transition-probability table, achieving competitive scheduling quality at negligible overhead. This selective application makes RIGEO practically deployable on resource-constrained fog controllers and scalable to the high task volumes characteristic of dense IoT deployments, where uniformly applying a meta-heuristic would saturate the scheduler before tasks can be dispatched.
Task scheduling faces multiple challenges, including energy efficiency, meeting deadline constraints, reducing latency, and ensuring reliability. The following subsections review relevant studies in these areas.
@@ I-A Energy Efficiency
Wang et al. [13] propose MEETS, a task scheduling algorithm for fog computing that determines optimal task allocation to fog nodes before scheduling tasks on each node. MEETS significantly reduces energy consumption while meeting all task deadlines; however, it assumes homogeneous fog networks and constant task arrival rates, which may not be realistic in practice.
Vispute et al. [12] propose a PSO-based task scheduling scheme that outperforms existing approaches in makespan, energy consumption, and execution time, though performance may vary across scenarios with different resource limitations.
@@ I-B Deadline Constraints
Azami et al. [1] incorporate two semi-greedy algorithms to minimize deadline violations for IoT tasks while ensuring QoS compliance, achieving a significant reduction in deadline violation time, though performance may vary under different real-world conditions.
@@ I-C Latency Reduction
Klatoun et al. [7] propose a task scheduling framework for IoMT that processes data closer to end-users, reducing latency and enhancing response times. However, the study focuses on specific scenarios, limiting broader applicability.
Hosseini et al. [6] present an algorithm using fuzzy logic and AHP to prioritize tasks based on completion time, energy, RAM, and deadlines. While effective at balancing cost and latency in mobile fog environments, its complexity could challenge resource-constrained deployments.
I-D Reliability and Fault Tolerance
Ghanavati et al. [5] propose a learning automata-based task scheduling approach that ensures reliable execution while optimizing response time and energy consumption, though its focus on specific scenarios may limit broader applicability.
@@ I-D Reliability and Fault Tolerance
Energy Deadline Latency Reliability
Efficiency Constraints Reduction
[13] 2018 × ×
[12] 2023 × × ×
@@ II-A Problem Statement
The problem is to find an optimal mapping $M:T\rightarrow F$ to achieve these objectives.
Each task $t_{i}$ is described by a tuple $(c_{i},m_{i},d_{i},a_{i},q_{i})$: CPU demand $c_{i}\in\{1,\ldots,5\}$ cycles, normalized memory demand $m_{i}\in[0,1]$, deadline $d_{i}\in[0,1]$, discrete arrival time-slot $a_{i}\in\{1,\ldots,T\}$ where $T=10$, and minimum required QoS level $q_{i}\in[0,1]$. Tasks arrive according to a uniform discrete model: each task independently draws its arrival slot $a_{i}\sim\mathcal{U}\{1,\ldots,T\}$, so batch sizes vary naturally across slots.
Each FN $f_{j}$ is described by a tuple $(C_{j},M_{j},\kappa_{j},\theta_{j},\varepsilon_{j},h_{j})$: CPU capacity $C_{j}\in[81,100]$ units, memory capacity $M_{j}\in[5,16]$ GB, cost rate $\kappa_{j}\in[0,1]$ per cycle, QoS throughput $\theta_{j}\in[0.5,1.5]$, energy consumption rate $\varepsilon_{j}\in[3,6]$ J/cycle, and energy harvest rate $h_{j}=0.003$. All FN parameters are drawn independently and uniformly from these ranges, yielding a heterogeneous network. The fog network is modeled as a fully connected graph: every FN pair $(f_{i},f_{j})$ shares a symmetric communication-preference weight $P(i,j)=P(j,i)\sim\mathcal{U}[0,1]$, which the RL agent refines at runtime.
The parameters used in the methodology are defined as follows:
@@ II-A3 Energy Consumption
II-B Simulation Environment
The simulation models a fog layer of $m=20$ heterogeneous FNs representing edge servers with diverse computational capabilities, as summarized in Table II. All FNs are fully interconnected (all-to-all mesh), and communication preferences between any pair of nodes are captured by the symmetric probability matrix $P$ described above. No FN represents a fixed real-world device; instead, the heterogeneous parameter distributions replicate the diversity found in practical deployments of commodity edge hardware.
Five task-load scenarios are evaluated (200, 300, 400, 500, and 600 tasks) over 50 independent runs with different random seeds for task generation and node initialization; reported values are averages across these runs. Tasks arrive over $T=10$ discrete time slots; within each slot, all arriving tasks are dispatched together as a batch. Meta-heuristic algorithms (IGEO and all baselines) are allocated $N_{pop}=50$ candidate solutions and $N_{iter}=2$ optimization iterations per batch. Traffic patterns for all 20 FNs are simulated over 200 historical time steps, with each node’s load drawn uniformly from $[1,100]$ units; FNs whose mean traffic exceeds the network-wide median are labeled high-traffic ($F_{\text{high}}$), the remainder low-traffic ($F_{\text{low}}$). The deadline threshold separating short- and long-deadline tasks is fixed at the 25th percentile of each incoming batch’s deadline distribution, and the traffic threshold is recomputed as the median at the start of each run.
Parameter Value
Fog nodes ($m$) 20
Network topology Full mesh (all-to-all)
Task loads evaluated 200, 300, 400, 500, 600
Time slots ($T$) 10
Independent runs 50
Population size ($N_{pop}$) 50
Iterations per batch ($N_{iter}$) 2
Traffic history length 200 steps
Max node traffic load 100 units
FN CPU capacity ($C_{j}$) $\mathcal{U}[81,100]$ units
FN memory ($M_{j}$) $\mathcal{U}[5,16]$ GB
FN energy rate ($\varepsilon_{j}$) $\mathcal{U}[3,6]$ J/cycle
FN energy harvest ($h_{j}$) $0.003$
FN cost rate ($\kappa_{j}$) $\mathcal{U}[0,1]$
FN QoS throughput ($\theta_{j}$) $\mathcal{U}[0.5,1.5]$
Deadline threshold 25th percentile of batch
Traffic threshold Median of node mean traffic
TABLE II: Simulation parameters
II-C Proposed Method (RIGEO)
@@
II-C1 Reinforcement Learning for Long-Deadline Tasks
RIGEO employs a table-based RL policy to schedule long-deadline tasks on $F_{\text{high}}$. The agent maintains the communication-preference matrix $P\in[0,1]^{m\times m}$ (initialized as described in Section II-B), where $P(s,f)$ encodes the learned preference for assigning a task to FN $f$ given that the preceding task was assigned to FN $s$. The matrix is updated online as each task in the batch is processed.
State. $s\in F_{\text{high}}$: the index of the most recently assigned high-traffic FN. The initial state is drawn uniformly at random from $F_{\text{high}}$.
Action. Given state $s$, the agent greedily selects the highest-valued reachable FN:
$a=\underset{f\in F_{\text{high}}}{\arg\max}\;P(s,f).$ (6)
Reward and penalty. After task $t_{i}$ is assigned to FN $a$, the response time $R_{i}=c_{i}/(C_{a}+1)$ is compared with deadline $d_{i}$. A met deadline rewards the transition (scales up by 10%); a violation penalises it (scales down by 10%):
$P(s,a)\leftarrow\begin{cases}\min\!\bigl(1,\;P(s,a)\times 1.1\bigr)&\text{if }R_{i}\leq d_{i},\\[2.0pt] \max\!\bigl(0,\;P(s,a)\times 0.9\bigr)&\text{if }R_{i}>d_{i}.\end{cases}$ (7)
State transition. The state advances to the assigned node, $s\leftarrow a$, so the policy adapts sequentially as tasks within the batch arrive.
This formulation limits per-task overhead to $\mathcal{O}(|F_{\text{high}}|)$, compared with $\mathcal{O}(N_{pop}\times N_{iter})$ for population-based methods, making it well-suited to long-deadline tasks whose relaxed timing constraint tolerates an incrementally improving policy rather than an exhaustive global search.
II-C2 Golden Eagle Optimization (GEO)
@@
II-C3 Improved Golden Eagle Optimization (IGEO)
@@
Negative step vectors in Eq. (9) indicate exploration tendency; positive vectors indicate exploitation. IGEO employs different genetic operators accordingly. For negative step vectors, mutation replaces addition in Eq. (10). Figure 2 illustrates the mutation operator, which randomly alters the parental genetic material.
@@
For positive step vectors, crossover operators provide efficiency. Figure 3 illustrates how genes from two parents exchange at predetermined cutting points. Eq. (12) shows the update for positive operations.
@@
(a) Energy Consumption (Joules)
(b) Deadline Violation Time (ms)
(c) Response Time (ms)
Figure 4: Performance metrics with varying task numbers (200-600) and 20 FNs over 50 runs: (a) Energy Consumption, (b) Deadline Violation, (c) Response Time
@@ III Results
Experimental results demonstrate that RIGEO reduces response time, minimizes deadline violation time, and optimizes energy consumption compared to other algorithms. Figure 4 shows the performance metrics across varying task numbers. The following sections examine RIGEO’s performance on each parameter.
@@ III-A Energy Consumption
Figure 4(a) illustrates the energy usage of FNs with varying task numbers (200 to 600) and 20 FNs. Energy consumption is calculated using Eq. (5).
@@ III-B Deadline Violation
RIGEO improves deadline adherence compared to other methods, as indicated by the decrease in violation times shown in Figure 4(b). The method consistently selects nodes with the shortest violation times, even in the presence of node shortages, demonstrating flexibility in addressing deadline breaches.
@@ III-C Response Time
Response time, the period required for the system response to a task, was systematically integrated into the fitness function as a critical parameter. Figure 4(c) illustrates RIGEO’s superior performance in minimizing response time compared to other approaches with 20 FNs.